OpenStudy (anonymous):

Coffee A and coffee B are mixed in the ratio x:y by weight. A costs $50/kg and B costs $40/kg. If the cost of A is increased by 10% while that of B is decreased by 15%, the cost of the mixture per kg remains unchanged. Find x:y.

OpenStudy (anonymous):

@Callisto @zzr0ck3r @hartnn @Fifciol

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Word problems are generally simple as far as solving goes, the tricky part is setting up what to solve... so let's see...

terenzreignz (terenzreignz):

Okay, so, coffee A costs 50 per kg, and there are x kilograms of coffee A... coffee B costs 40 per kg, and there are y kilograms of coffee B... so we have 50x + 40y = k Where k is the cost...

terenzreignz (terenzreignz):

k is the cost of the mixture.... catch me so far?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, if we increase the cost of x by 10%... what is its new price per kg?

terenzreignz (terenzreignz):

lol... @kryton1212 you better make your presence felt :P

OpenStudy (anonymous):

1.1x ...

terenzreignz (terenzreignz):

okay, so what is 1.1*50?

OpenStudy (anonymous):

55...

terenzreignz (terenzreignz):

That's correct, so that's the new price of coffee A. What about coffee B? We *decrease* its cost by 15% so what is its new price?

OpenStudy (anonymous):

55x+34y=k?

terenzreignz (terenzreignz):

Yes, that is correct :) And since we have 50x + 40y = k and 55x + 34y = k then clearly 50x + 40y = 55x + 34y

terenzreignz (terenzreignz):

Can you take it from there? :)

OpenStudy (anonymous):

yes, 5x=6y

OpenStudy (anonymous):

6:5

terenzreignz (terenzreignz):

Good job ^_^

OpenStudy (anonymous):

thanks:))